OpenStudy (anonymous):

Suppose a triangle has two sides of length 3 and 4 and that the angle between these two sides is pi/3. What is the length of the third side of the triangle?

OpenStudy (imstuck):

Are we to assume this is not a right triangle?

OpenStudy (anonymous):

yes i think so

OpenStudy (imstuck):

Are you studying the Law of Sines, by any chance?

OpenStudy (imstuck):

Law of Cosines, perhaps?

OpenStudy (imstuck):

You would need to use the Law of Cosines for this to solve for the missing side.

OpenStudy (imstuck):

If this was your triangle (NOT a right triangle):

OpenStudy (imstuck):

You would need the Law of Cosines to find side a, your missing side. The Law of Cosines for this set up would be this:\[a ^{2}=b ^{2}+c ^{2}-2bc(cosA)\]

OpenStudy (imstuck):

Are you even here anymore? Am I talking to no one?

OpenStudy (anonymous):

no i'm here

OpenStudy (anonymous):

thank you by the way :)

OpenStudy (imstuck):

Cool. Do you know the Law of Cosines?

OpenStudy (anonymous):

not well but if i have the equation i can plug things in

OpenStudy (imstuck):

Notice, also, that I converted the pi/3 to a degree measure of 60. I don't know about you, but my brain does not operate in radians. Degrees we are all familiar with. So always try to convert to degrees if you can, to make things easier.

OpenStudy (anonymous):

alrighty

OpenStudy (imstuck):

Now that you have the equation, you can plug in your numbers. Try it. Let me see what you come up with. Show me your answer, ok?

OpenStudy (anonymous):

alright gimme a second

OpenStudy (imstuck):

Make sure your calculator is in degrees for this, not radians.

OpenStudy (anonymous):

a2=3^2+4^2−2(3)(4)(cos60)

OpenStudy (anonymous):

that would be just plugging everything in right?

OpenStudy (imstuck):

Go on....

OpenStudy (imstuck):

BRB...just in case you post and I do not respond immediately.

OpenStudy (anonymous):

a^2=9+16-24(.799)

OpenStudy (imstuck):

Ok, the cosine of 60 is actually .5 I am not sure where you got .799

OpenStudy (imstuck):

But everything else is good.

OpenStudy (imstuck):

How did you get .799?

OpenStudy (anonymous):

my calculator?

OpenStudy (anonymous):

oh shoot i typed the wrong number my bad

OpenStudy (imstuck):

Try it again. What do you get this time?

OpenStudy (anonymous):

.5

OpenStudy (imstuck):

Cool! NOW do the calculations and tell me what the length of side a is, please! ; )

OpenStudy (anonymous):

sqrt a=sqrt .5, am i doing this right?

OpenStudy (imstuck):

Put a wrong sign in there...let me retype...

OpenStudy (imstuck):

\[a ^{2}=16+9-24(.5)\]\[a ^{2}=25-12\]\[a ^{2}=13\]\[a=\sqrt{13}\]

OpenStudy (imstuck):

See that?

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

yeah i did 25-24 and then multiplied by .5

OpenStudy (anonymous):

thank you :)

OpenStudy (imstuck):

Watch your order of operations here. When you do these, always always do the multiplication first, then subtraction.

OpenStudy (anonymous):

could you help me with one more questiom? yeah I always forget that part and it messes me up.

OpenStudy (imstuck):

I would be happy to help!

OpenStudy (imstuck):

BRB once again....just give me a sec

OpenStudy (anonymous):

I'm just drawing the triangle

OpenStudy (anonymous):